OpenStudy (ammarah):

Factor the sum or difference of cubes. y^3-8

OpenStudy (ranga):

\[\Large a ^{3} - b ^{3} = (a - b)(a ^{2} + ab + b ^{2})\]

OpenStudy (ammarah):

is that the final answer?

OpenStudy (ammarah):

i dont know how to find b

OpenStudy (ranga):

No that is the identity you will use to factor y^8 - 8 Write y^3 - 8 as y^3 - 2^3 and compare it to the identity. a = y and b = 2

OpenStudy (ammarah):

so is it y-2 ....i dont know the second part

OpenStudy (ranga):

Put a = y and b = 2 on the right side of the formula I gave above.

OpenStudy (ammarah):

(y-2)(y^2+2y+4)?

OpenStudy (ammarah):

did i get it right???^^

OpenStudy (ranga):

Yes.

OpenStudy (ammarah):

what do i do for this one: 64n^3-27

OpenStudy (ammarah):

do i factor it first?

OpenStudy (ranga):

No. Use the fact that 64 is cube of 4. 27 is cube of 3.

OpenStudy (ammarah):

so a = 4 and b=3?

OpenStudy (ranga):

64n^3-27 = (4n)^3 - 3^3 a = 4n and b = 3

OpenStudy (ammarah):

wait so do i plug it into the same equation above???

OpenStudy (ranga):

Yes.

OpenStudy (ammarah):

I got (4n-3)(4n^2_12n+9)

OpenStudy (ranga):

No. a = 4n, b = 3. A62 = (4n)^2 = 16n^2, ab = +12n, b^2 = +9

OpenStudy (ammarah):

? i dont get it

OpenStudy (ammarah):

whats A62???

OpenStudy (ranga):

This is a typo. It should be a^2 a^2 = (4n)^2 = 16n^2, ab = +12n, b^2 = +9

OpenStudy (ammarah):

can u please write the answer....ill get it when it is shown...

OpenStudy (ranga):

64n^3-27 = (4n)^3 - 3^3 = (4n - 3)(16n^2 + 12n + 9)

OpenStudy (ammarah):

ok how about this one: 2w^3+54

OpenStudy (ammarah):

a=2 b=?

OpenStudy (ranga):

Factor the 2 out first. 2w^3+54 = 2(w^3 + 27) = 2(w^3 + 3^3) Here you have to use a different formula. Previously it was difference of two cubes. Here it is sum of two cubes:\[\Large a ^{3} + b ^{3} = (a + b)(a ^{2} - ab + b ^{2})\]

OpenStudy (ammarah):

what does a and b equal?

OpenStudy (ranga):

compare a^3 + b^3 to w^3 + 3^3. You should be able to identify a and b having done two problems already.

OpenStudy (ammarah):

2 and 3?

OpenStudy (ammarah):

is that it??????

OpenStudy (ranga):

I am sorry. You want answers. But having done two problems you should be able to identify a and b and put it in the formula. I can't keep giving answers for the same problem with just slightly different numbers.